OpenStudy (anonymous):

X^2+5X+6/X^2-6X+5 x X^2+X-30/X^2+9X+18

OpenStudy (anonymous):

Rational Expressions...Can I get a step by step procedure to solve this please.

OpenStudy (kropot72):

All the numerators and denominators can be factorised. This is the first step. For example can you factorise the following: \[x ^{2}+5x+6=(x+?)(x+?)\]

OpenStudy (anonymous):

x2+5x+6=(x+3)(x+2)

OpenStudy (kropot72):

Good! So can you factorise the other terms?

OpenStudy (anonymous):

so after i do that, what is the next step?

OpenStudy (kropot72):

You will find that six of the factors will cancel and you will be left with an expression of the form: \[\frac{x+?}{x-?}\]

OpenStudy (anonymous):

Oh wow, I feel so stupid haha. Thanks. I did a very similar problem yesterday and the method just slipped out of my mind.

OpenStudy (kropot72):

You're welcome :)